OpenStudy (anonymous):

5f-(3f+4)=20 The parantheses stand for absolute value

OpenStudy (anonymous):

solve for f

OpenStudy (anonymous):

can u please help

OpenStudy (anonymous):

|3f+4| = 5f-20 so we know |x| = a then \[x= \pm a\]

OpenStudy (anonymous):

-1*abs(3f+4)=20-3f abs(3f+4)=-20+3f

OpenStudy (anonymous):

what does abs mean?

OpenStudy (anonymous):

absolute value

OpenStudy (anonymous):

oh.. so u pretty much distribute it like if it were just a regular parantheses?

OpenStudy (anonymous):

no

OpenStudy (jdoe0001):

so \(\bf 5f-|3f+4|=20\) keep in mind that |whatever| really means +1(whatever) AND -1(whatever) thus you'd get 2 values for "x", then \(\bf 5f-|3f+4|=20\\ \implies 5f-20 = |3f+4|\\ 5f-20 = +1(3f+4) \implies 5f-20 = 3f+4\\ 5f-20 = -1(3f+4) \implies 5f-20 = -3f-4\)

OpenStudy (anonymous):

im confused on how you reached the third step?

OpenStudy (anonymous):

sorry i meant the second step

OpenStudy (jdoe0001):

hmmm which 2nd step?

OpenStudy (anonymous):

hurry please i g2g

OpenStudy (anonymous):

5f-20= abs 3f +4

OpenStudy (anonymous):

-1*abs(3f+4)=20-3f subtracted 3f from both sides to get it on the right NOTICE: -abs(3f+4) and -1*abs(3f+4) are same but it's just easier to see that it is multiplied by a negative. It is just shorthand. we multiply both sides by -1 -1*-1*abs(3f+4) = -1*(-20+3f) abs(3f+4)=-20+3f Now as jdoe says x=±a same as 3f+4= (-20+3f) and 3f+4= -(-20+3f) I think that is how it works you have to solve both equations but now you got rid of absolute value sign :)

OpenStudy (jdoe0001):

\(\bf 5f-|3f+4|=20 \implies 5f-20 = |3f+4| \ \ ?\)

OpenStudy (anonymous):

yes that one

OpenStudy (jdoe0001):

ohh, all I did was isolate the absolute expression, so I can do the +1/-1 part with less clutter

OpenStudy (anonymous):

oh ok so the two answers are both the same just negative and positive?

OpenStudy (jdoe0001):

yes, absolute value expressions give you 2 values

OpenStudy (jdoe0001):

so x will have 2 values

OpenStudy (anonymous):

so this is for every absolute value equation? same solution just negative and positive

OpenStudy (jdoe0001):

the expression turns negative and positive, yes

OpenStudy (anonymous):

alright thanks ;D how do I give a medal?

OpenStudy (jdoe0001):

heheh, no need heh

OpenStudy (anonymous):

lol i figured u dont even have a profile pic. thanks again

OpenStudy (jdoe0001):

well, yes I was going to upload this pic of mine http://www.fabioinc.com/FABIO_MODEL%20folder/Fabio_M_blueShirt_gap_web.jpg but I figure it might create a riot and a lawsuit from gap inc. so tis ok =)

OpenStudy (jdoe0001):

ok ok, so it's not my pic =)